Grade 12Electrostatics

A small block of mass m, charge q is kept at the top of a smooth inclined plane of angle 30 placed in an elevator moving upward with an acceleration `a`. Electric field E exists between the vertical side walls of the elevator . The time taken by the block to come to the lowest point of inclined plane is?

Shalini sai
8 Years agoGrade 12

1 Answer

Rituraj Tiwari
5 Years ago

To determine how long it takes for a small block of mass \( m \) and charge \( q \) to slide down an inclined plane while the elevator is accelerating upward, we need to analyze the forces acting on the block. This scenario combines gravitational, electric, and inertial forces due to the elevator's acceleration. Let's break this down step by step.

Analyzing Forces at Play

First, let’s identify the forces acting on the block:

  • Gravitational Force: The weight of the block acts downward, which is given by \( mg \), where \( g \) is the acceleration due to gravity.
  • Normal Force: The normal force acts perpendicular to the inclined plane.
  • Electric Force: The electric field \( E \) exerts a force \( F_e = qE \) on the block, which can either aid or oppose its motion depending on the direction of \( E \).
  • Inertial Force: Since the elevator is accelerating upward with acceleration \( a \), we consider an effective gravitational acceleration of \( g' = g + a \) acting downward on the block.

Resultant Acceleration Down the Incline

Now, let's express the forces in terms of their components along the incline. The gravitational force component acting down the incline is:

\( F_{gravity} = mg \sin(\theta) = mg \sin(30^\circ) = \frac{mg}{2}

The effective gravitational force due to the elevator's upward acceleration also contributes to the net force. So, we can express the net force acting on the block down the incline as:

\( F_{net} = mg \sin(30^\circ) + qE = \frac{mg}{2} + qE

Now, applying Newton's second law, we can relate this force to the acceleration:

\( ma = F_{net} \)

Thus, the acceleration \( a_{incline} \) of the block along the incline becomes:

\( ma_{incline} = \frac{mg}{2} + qE

From this, we can solve for \( a_{incline} \):

\( a_{incline} = \frac{g}{2} + \frac{qE}{m}

Time to Slide Down the Inclined Plane

Next, we need to find the time it takes for the block to slide down the incline. Let’s denote the length of the incline as \( L \). Using the equation of motion for uniformly accelerated motion:

\( L = \frac{1}{2} a_{incline} t^2

We can rearrange this to solve for time \( t \):

\( t^2 = \frac{2L}{a_{incline}}

Substituting for \( a_{incline} \) gives:

\( t^2 = \frac{2L}{\frac{g}{2} + \frac{qE}{m}}

Finally, taking the square root, we arrive at the expression for the time taken by the block to reach the lowest point on the inclined plane:

\( t = \sqrt{\frac{2L}{\frac{g}{2} + \frac{qE}{m}}}

Final Thoughts

This formula effectively captures the interplay of gravitational, electric, and inertial forces acting on the block. The time taken for the block to slide down the incline depends on the incline length \( L \), the mass \( m \), the charge \( q \), the electric field \( E \), and the gravitational acceleration \( g \). By plugging in the appropriate values, you can find the time for any specific scenario. If you have specific values for these variables, we can calculate the numerical time together!